Shaypop:

The quadrilateral shown is a rhombus. What is m ‹ 2?

snowflake0531:

pic

Shaypop:

AZ:

You have to know how the angles are related to each other

1 attachment
AZ:

so angle 2 = 32\(^o\) as well and so does angle 4 and 5

AZ:

to find angle 2, you can remember that the two adjacent angles add up to 180 so angles 1, 2, 4, and 5 add up to 180 and we know that angles 4 and 5 are 32 degrees so m<1 + m<2 + 32\(^o\) + 32\(^o\) = 180\(^o\) and remember that angle 1 and angle 2 are equal to each and they're also equal to angle 6 and 7 so if angle 2 is x degrees, then angle 1 is also x degrees so you have x + x + 32 + 32 = 180

AZ:

solve for x, and you're all good to go

Shaypop:

64

Shaypop:

32+32

AZ:

yes and subtract that on both sides then

Shaypop:

180-64

AZ:

good!

AZ:

don't forget that x + x = 2x so now you have 2x = 180 - 64

Shaypop:

58

AZ:

Perfect! Well done :)

Shaypop:

Thank u

AZ:

You're most welcome :)